OpenStudy (kobeni-chan):

(1/4) - (1/x) = 1/(x+15) Solve for x by making a common denominator

OpenStudy (michele_laino):

please, note that common denominator is 4*x*(x+15)

OpenStudy (kobeni-chan):

do I distribute 4 and x into the parentheses?

OpenStudy (michele_laino):

No I don't think. I think you have perform this calculus: \[\frac{... }{ 4*x*(x+15) }-\frac{ ... }{ 4*x*(x+15) }=\frac{ ... }{ 4*x*(x+15) }\]

OpenStudy (michele_laino):

now I ask you what is [4*x*(x+15)]/4?

OpenStudy (michele_laino):

@kobeni-chan

OpenStudy (kobeni-chan):

x(x+15) right?

OpenStudy (michele_laino):

that's right, that is your first numerator, namely the leftmost numerator

OpenStudy (michele_laino):

now I ask you what is [4*x*(x+15)]/x?

OpenStudy (kobeni-chan):

4(x+15) right? I don't think I'm supposed to take x out of the parentheses

OpenStudy (michele_laino):

that's right! that is your second numerator, namely at middle position: